# When Text Embedding Meets Large Language Model: A Comprehensive Survey

### v3

@@ 4.3.3 Supervised Contrastive Learning (SCL)
Prompt-Based Learning
@@
Prompt-based learning [218] aligns the fine-tuning form of downstream tasks with that of the pre-training task, enabling the mode learned during pre-training to be fully utilized. Some prompt-based training-free methods, such as PromptEOL [170], PromptReps [184] and Echo [190], Further, BeLLM [178] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
Instruction Tuning
General-purpose embedding models confront inherent challenges stemming from the diverse and often conflicting objectives across various downstream tasks. For instance, a sentence pair deemed semantically similar may prove irrelevant in a retrieval context if it fails to address a specific query. Early approaches attempted to mitigate this issue by differentiating symmetric and asymmetric retrieval through the use of prefixes, such as ’query:’ and ’passage:’. However, with the expanding landscape of applications, including classification, clustering, retrieval, and semantic similarity assessment, instruction-based prompting has emerged as a promising strategy. This approach aims to enhance task-specific performance by explicitly incorporating task descriptions within the input.
@@
Instructor [67] firstly pioneered instruction tuning for embedding models, building upon GTR [179] and curating a dataset encompassing various tasks. Building upon this foundation, Data-CUBE [219] advanced the field by implementing curriculum learning within multitask instruction tuning. TART [51] specifically targeted retrieval tasks, constructing the BERRI dataset, which covers 40 distinct retrieval scenarios. The robust instruction-following capabilities of large language models (LLMs) have catalyzed the development of LLM-based embedding models with instruction tuning. InBedder [197] proposes a novel methodology for training instruction-following sentence embedders by exploiting the generative capacity of LLMs. Rather than directly manipulating embedding vectors, InBedder fine-tunes LLMs on abstractive question answering, extracting sentence embeddings from the averaged hidden states of generated answers during inference. E5-Mistral [157] utilizes a large, LLM-generated, instruction-rich training dataset to tune LLM embeddings. Notably, this model initially applied instructions only 0to the query side, a practice subsequently adopted by numerous subsequent works [182, 183, 153, 187, 188, 168, 169, 198, 199, 200, 174, 202, 203]."
In-Context Tuning
@@
The practice of in-context tuning [220], which augments training inputs with contextual examples, has proven effective in enhancing a model’s understanding of contextual dependencies. Following this trend, recent research [187, 202] has explored the adaptation of in-context tuning for LLM-based embedding models. These approaches involve fine-tuning pre-trained LLMs with in-context examples that share semantic similarity with the target input, resulting in improved performance on various downstream tasks."
4.3.4 Multi-Task Learning
@@
Beyond the standard contrastive learning objective, recent research explores the integration of auxiliary learning objectives during supervised training to enhance representation quality and preserve generative capacities in Large Language Models (LLMs). This multi-objective approach aims to broaden the applicability of LLMs across diverse downstream tasks.
Piccolo2 [221] adopts task-specific loss functions, employing contrastive loss (CL) for retrieval and margin loss for semantic textual similarity (STS). This task-adaptive loss strategy optimizes performance for distinct modalities. Similarly, Contextual Document Embeddings (CDE) [222] leverages neighboring document information to generate contextualized embeddings, significantly improving neural retrieval, particularly in out-of-domain scenarios, compared to traditional bi-encoders.
GritLM [169] demonstrates that simultaneous training with contrastive learning and Next Token Prediction (NTP) enables LLMs to achieve robust text representation while maintaining their native generative abilities. This dual-objective approach balances representational and generative demands. ChatRetriever [183] further refines LLMs for retrieval by integrating contrastive learning with masked instruction tuning on high-quality conversational data, thereby enhancing complex session understanding.
@@
DEBATER [188] introduces a continuous thinking process to augment dense retrieval. This method generates a sequence of tokens representing a chain of thought prior to producing the final query representation. Document similarity is then determined by the maximum similarity score between any token in the chain and the document. Training incorporates both contrastive learning and self-distillation to refine the representation.
ULLME [201] proposes Generation-Representation Learning (GRL), a fine-tuning technique that jointly optimizes contrastive learning and generation objectives. GRL minimizes the Kullback-Leibler (KL) divergence between similarity scores derived from learned representations and probability distributions from generation, ensuring internal consistency between the representational and generative aspects of the model.
O1-Embedder [189] utilizes a single LLM for both query expansion and text embedding. The query expansion is modelled as a next token prediction task, where a large-size LLM is used to generate the supervised queries. At inference time, the embedding of the original query and O1-Embedder’s newly generated query are pooled equally as the final query embedding.
4.3.5 Multi-Stage Training
@@
Neither BERT nor current LLMs are primarily designed as embedding models during their pre-training phases. Consequently, there has been considerable research focused on adapting the base pre-trained language models before supervised training to make them more suitable for embedding tasks. During the BERT pre-training era, representative works include CoCondenser [223] and RetroMAE [224], which have designed distinct training objectives to enhance BERT’s text embedding capabilities. In the era of LLMs, similar approaches have been extended to improve LLMs’ embedding abilities. For example, Llama2Vec [180] employs various templates that enable the model’s hidden states to predict which words from the original sentence and the subsequent sentence are included. This method serves as a secondary pre-training technique to enhance the model’s capacity to capture global semantics with the hidden state from a position. Llama2Vec utilizes two auxiliary pre-training tasks, EBAE (Embedding-Based Auto-Encoding) and EBAR (Embedding-Based Auto-Regression), to adapt the foundation model, substantially improving the model’s fine-tuned performance on a variety of dense retrieval benchmarks.
@@
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [179], GTE-Qwen2 [198], and BMRetriever [182] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage provided by typically over a billion data points significantly improves the model’s domain generalization capabilities. Additionally, NV-Retriever proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). Specifically, NV-Retriever [153] uses only retrieval-supervised data with in-batch negatives alongside mined hard negatives for the first stage while blending data for retrieval tasks with datasets from other tasks for the second stage.
@@ 4.4 Commercial Service
Many AI companies provide generic text embedding services to support traditional NLP tasks and RAG. In English text embedding service, it mainly includes OpenAI  [^11] https://openai.com/index/introducing-text-and-code-embeddings, Google  [^12] https://cloud.google.com/vertex-ai/generative-ai/docs/embeddings/get-text-embeddings, Amazon  [^13] https://docs.aws.amazon.com/bedrock/latest/userguide/titan-embedding-models.html, Alibaba Cloud [^14] https://www.alibabacloud.com/help/en/model-studio/developer-reference/general-text-embedding, Voyage  [^15] https://docs.voyageai.com/docs/embeddings and Cohere  [^16] https://cohere.com/blog/introducing-embed-v3. Due to the trade secrets involved, only a few companies have revealed some of the technology behind their services. For example, OpenAI simultaneously provides text embeddings with different dimensional settings for different levels of time-storage sensitivity in practice, where matryoshka representation learning (MRL) [225] is considered to be the key to obtaining the embedding with flexible dimensions using only one encoder. Google demonstrates that its Gecko [156], which achieves superior performance on 1B scale and 786 dimensions, relies heavily on a two-stage synthetic data generation process with LLMs: the first step generates diverse tasks and queries based on the instruction and the second step generates positive and negative samples based on the obtained tasks and queries. In addition, there have been methods [226] that attempt to distill knowledge through embeddings obtained from commercial APIs, obtaining a local embedder with similar performance successfully.

### v4

@@ 4.3.3 Supervised Contrastive Learning (SCL)
Prompt Tuning (Prompt)
@@
Prompt Tuning [239] aligns the fine-tuning form of downstream tasks with that of the pre-training task, enabling the mode learned during pre-training to be fully utilized. Some prompt-based training-free methods, such as PromptEOL [169], PromptReps [181], and Echo [189], further, BeLLM [173] explores the impact of backward dependencies in LLMs and then proposes a backward dependency enhanced large language model (BeLLM) that transforms certain attention layers from unidirectional to bidirectional. This modification allows the model to learn more nuanced sentence embeddings, improving performance across various STS tasks and demonstrating the value of incorporating backward dependencies.
Instruction Tuning (Instruction)
@@
General-purpose embedding models confront inherent challenges stemming from the diverse and often conflicting objectives across various downstream tasks. For instance, a sentence pair deemed semantically similar may prove irrelevant in a retrieval context if it fails to address a specific query. Early approaches attempted to mitigate this issue by differentiating symmetric and asymmetric retrieval through the use of prefixes, such as ’query:’ and ’passage:’. However, with the expanding landscape of applications, including classification, clustering, retrieval, and semantic similarity assessment, instruction-based prompting has emerged as a promising strategy. This approach aims to enhance task-specific performance by explicitly incorporating task descriptions within the input. Instructor [67] firstly pioneered instruction tuning for embedding models, building upon GTR [175] and curating a dataset encompassing various tasks. Building upon this foundation, Data-CUBE [240] advanced the field by implementing curriculum learning within multitask instruction tuning. TART [51] specifically targeted retrieval tasks, constructing the BERRI dataset, which covers 40 distinct retrieval scenarios. The robust instruction-following capabilities of large language models (LLMs) have catalyzed the development of LLM-based embedding models with instruction tuning. InBedder [196] proposes a novel methodology for training instruction-following sentence embedders by exploiting the generative capacity of LLMs. Rather than directly manipulating embedding vectors, InBedder fine-tunes LLMs on abstractive question answering, extracting sentence embeddings from the averaged hidden states of generated answers during inference. E5-Mistral [157] utilizes a large, LLM-generated, instruction-rich training dataset to tune LLM embeddings. Notably, this model initially applied instructions only 0to the query side, a practice subsequently adopted by numerous subsequent works [179, 180, 153, 184, 185, 193, 197, 218, 217, 198, 200, 214, 212].
In-Context Tuning (Example)
@@
The practice of in-context tuning [241], which augments training inputs with contextual examples, has proven effective in enhancing a model’s understanding of contextual dependencies. Following this trend, RARe [184] and BGE-ICL [214] have explored the adaptation of in-context tuning for LLM-based embedding models. These approaches involve fine-tuning pre-trained LLMs with in-context examples that share semantic similarity with the target input, resulting in improved performance on various downstream tasks.
4.3.4 Next Token Prediction
Next Token Prediction is a primary paradigm in the pre-training and supervised fine-tuning of LLMs. Obviously, the supervisory signal for next token prediction resides in the vocabulary space, inherently resulting in a dimensionality gap with the embedding space. This makes it challenging to design supervision signals in the vocabulary space and correctly influence the quality of text embeddings. Although theoretical research in this area remains scarce, interestingly, several successful empirical cases have demonstrated the potential of this generative task in text embedding. Note that the works introduced in this section replace contrastive learning with next token prediction as the primary task. For the work that regards next token prediction as a pretext task before contrastive learning, please refer to Section 4.3.7.
Unsupervised Next Token Prediction (UNTP)
Text2Token [208] is inspired by empirical findings [232]: when obtaining a text embedding from the LLM-based embedder, the tokens with the highest decoding probability are the key tokens of the input text. Therefore, Text2Token explores whether the opposite approach holds: if leveraging a given target token distribution as the supervisory signal, the embedders can learning the embedding similar to contrastive fine-tuning. Using data-based statistical methods or model-based filtering methods, Text2Token constructs the target distributions for key tokens in an unsupervised manner and achieves performance comparable to unsupervised LLM2Vec [193].
Supervised Next Token Prediction (SNTP)
Inbedder [196] explores training models using query-answer pairs to enable the LLM-based embedders to follow instructions. The authors argue that for specific questions about textual content, the consistency of LLMs’ responses reflects the semantic similarity between texts on the subject being inquired. At the same time, to ensure LLMs provide immediate and relevant responses to questions, it is necessary to guide them in generating short yet informative replies. Given these considerations, the author employed plenty of short-answer question-answer pairs to train LLMs and empirically demonstrated that last pooling (corresponding to the generation of the first word) yields the best embedding performance.
4.3.5 Reinforcement Learning (RL)
Reinforcement Learning from Human Feedback (RLHF) [167] serves as the final step in current advanced LLM training pipelines, designed to align with human preferences. The reinforcement learning algorithms within RLHF have evolved from the earliest Proximal Policy Optimization (PPO) [242] to methods such as Direct Preference Optimization (DPO) [243] and Group Relative Policy Optimization (GRPO) [244]. Among these algorithms, GRPO dispenses with the need for a value model in PPO and high-quality preference data in DPO, making it the primary technical approach currently explored within the text embedding community [187, 204].
The key to GRPO lies in sampling a group of candidate solutions for each sample and conducting group evaluations to estimates the group-relative advantage of each solution. Therefore, unlike previous methods, RL-based approaches require randomness in the embedding process to achieve sampling. The most intuitive approach is therefore to incorporate the generated content into the pooling strategy. The core difference of these approaches lies in designing a reward function, and We highlight the distinctions in reward design as below.
Search-R3 [187] focuses on retrieval tasks, with its reward simultaneously considering (1) format correctness and (2) the Discounted Cumulative Gain (DCG) metric. The first term ensures the embeddings’ accessibility: the text embeddings is originated from a special token generated after the reasoning content, thus the reward remains non-zero only when the generated content contains this special token. And the second term directly uses scaled DCG as the reward metric—a straightforward approach. However, since the embedding space continuously evolves during training, re-encoding the entire corpus is impractical. Search-R3 addresses this by introducing a specialized RL environment that incorporates sample forms, document-simplified embeddings, and localized graph refreshes. This environment updates embeddings in only a portion of the corpus at each iteration, mitigating computational overhead.
GRACE [204] focuses on general embedding, and its rewards draw upon the historical successes of (1) contrastive learning and (2) hard-to-learn example mining, while also incorporating the (3) consistency of GPRO sampling. Specifically, the first term emphasizes that the similarity between anchors and positive examples should be higher than that with negative examples; the second term encourages minimizing the similarity to hard negative examples from other samples; while the third term constrains the consistency among embeddings obtained through multiple sampling solutions. Note that the latter two items do not require additional annotation. For the first item, GRACE provides an unsupervised version that encourages the embedding similarity between the two anchor’s sampling solutions to be as high as possible. Interestingly, the embedder trained through this algorithm did not lose its generation capabilities.
[…]
@@ 4.3.5 Reinforcement Learning (RL)
Beyond the standard contrastive learning objective, recent research explores the integration of auxiliary learning objectives during supervised training to enhance representation quality and preserve generative capacities in LLMs. This multi-objective approach aims to broaden the applicability of LLMs across diverse downstream tasks.
Hybrid Loss for Different Data Format
Contrastive learning, which merely classifies samples as positive or negative examples, often loses finer-grained information such as similarity scores of STS datasets. Advanced models, however, require the full utilization of annotations across various tasks to achieve optimal performance [245]. To fully leverage the fine-grained annotations from the STS task, CoDiEmb [209] simultaneously introduced a Pearson Loss [246] to fit similarity scores, a newly proposed Rank KL-divergence Loss to fit similarity rankings, and a RL algorithm called Preference Rank Optimization (PRO) [247] to penalize text pairs with reversed rankings. QZhou-Embedding [213] uses CoSENT Loss [248] to reinforce the ranking relationship between sample similarity scores. Search-R3 [187] introduces Triplet Loss [249] to widen the similarity gap between anchor-positive and anchor-negative pairs.
Contrastive Learning + Generative Objective
Relying solely on contrastive learning can cause LLMs to lose nearly all their generative capabilities [197]. To enable a model to possess both generative and embedding capabilities simultaneously, or to enhance embedding capabilities by generating additional reasoning steps, it is necessary to find ways to preserve the generative capabilities of LLMs during training. A straightforward approach is to combine generative tasks with contrastive learning for joint training. For example, GritLM [197] demonstrates that simultaneous training with contrastive learning and Next Token Prediction (NTP) enables LLMs to achieve robust text representation while maintaining their native generative abilities. This dual-objective approach balances representational and generative demands. ChatRetriever [180] further refines LLMs for retrieval by integrating contrastive learning with masked instruction tuning on high-quality conversational data, thereby enhancing complex session understanding. O1-Embedder [186] utilizes a single LLM for both query expansion and text embedding. The query expansion is modeled as a next token prediction task, where a large-scale LLM is used to generate the supervised queries. At inference time, the embedding of the original query and O1-Embedder’s newly generated query are pooled equally as the final query embedding. ULLME [199] proposes Generation-Representation Learning (GRL), a fine-tuning technique that jointly optimizes contrastive learning and generation objectives. GRL minimizes the Kullback-Leibler (KL) divergence between similarity scores derived from learned representations and the generation probability distributions, ensuring internal consistency between the representational and generative aspects of the model.
Contrastive Learning + Self-Distillation
@@
DEBATER [185] introduces a continuous thinking process to augment dense retrieval, which generates a special token sequence representing the chain of thought before producing the final document embedding. During training, the maximum similarity score between any token in the chain and the query then determines the final similarity. And an additional self-distillation term that preserves the use of only the last token’s embedding can achieve rankings as close as possible to those from all tokens. Therefore, DEBATER can use only the embedding of the last token during inference, reducing computational complexity.
Contrastive Learning + Regularization
Mistral-SPLADE [183] use the FLOPs regularization [250], applying to the representations in the vocabulary space to ensure the desired sparsity factor. GIRCSE [206] leverages the generative capabilities of LLMs to produce multiple soft tokens, each weighted by multiple token embeddings. The last hidden state of all soft tokens is averaged to serve as the embedding for the entire text. Beyond contrastive learning, the authors introduce Iterative Contrastive Refinement (ICR) to constrain the contrastive loss, ensuring that the loss calculated from the average pooling of the $K+1$ soft tokens is lower than that calculated from the $K$ soft tokens.
[…]
@@
Neither BERT nor current LLMs are primarily designed as embedding models during their pre-training phases. Consequently, considerable research focuses on adapting the base pre-trained language models before contrastive learning to make them more suitable for embedding tasks. During the BERT pre-training era, representative works include CoCondenser [251] and RetroMAE [252], which have designed distinct training objectives to enhance BERT’s text embedding capabilities. Shifting to the era of LLMs, the lack of semantic aggregation capability in language models remains unchanged; therefore, similar approaches have been extended to improve LLMs’ embedding abilities. For the unsupervised setting, Llama2Vec [177] employs various templates that enable the model’s hidden states to predict which words from the original sentence and the subsequent sentence are included. Correspondingly, it utilizes two auxiliary pre-training tasks, EBAE (Embedding-Based Auto-Encoding) and EBAR (Embedding-Based Auto-Regression). These two tasks serve as a secondary pre-training technique to enhance the model’s ability to capture global semantics using the hidden state from a given position. For the supervised setting, Anchor [207] introduces the bi-directional reconstruction for query-document pairs. Specifically, the authors introduce Embedding-Based Query-to-Document (EBQ2D) to reconstruct document text from query embedding and Embedding-Based Document-to-Query (EBD2Q) to reconstruct query text from document embeddings. Following the long text compression method, AutoCompressor [253], AutoRegEmbed [174], append multiple special tokens to the end of the query text, and the document is predicted by extracting the last hidden state corresponding to these special tokens. Some pretext tasks are highly relevant to the design of other components, such as mitigating the impact of bidirectional attention switching on LLMs [193] or ensuring the embedder maintains format correctness during inference [187]. These methods have been introduced in detail in other sections, so we will not elaborate further here.
@@
During the supervised learning phase, numerous studies have explored multi-stage contrastive learning training methods to enhance further the generalization and versatility of the final embedding models. For instance, influenced by earlier work from the BERT era, models like GTR [175], GTE-Qwen2 [218], and BMRetriever [179] have investigated two-stage training strategies that combine weakly supervised contrastive learning (WCL) with supervised contrastive learning (SCL). Weakly supervised contrastive learning primarily leverages a large amount of weakly supervised relevance data collected from public domains (e.g., neighboring text spans and question-answer pairs from the QA community) for training. Although this data may contain noise, the extensive domain coverage, typically comprising over a billion data points, significantly improves the model’s domain generalization capabilities. Additionally, NV-Retriever proposes a phased supervised training approach tailored to different tasks (e.g., retrieval and STS tasks). Specifically, NV-Retriever [153] utilizes only retrieval-supervised data with in-batch negatives, alongside mined hard negatives, for the first stage, while blending data for retrieval tasks with datasets from other tasks for the second stage.
4.3.8 Model Merging (MG)
Model Merging [254, 255] was initially proposed to average multiple checkpoints fine-tuned on the same task, enhancing the model’s generalization performance. Following model soups approach [255] in this line, Gemini Embedding [216] weights multiple final checkpoints obtained from different runs to enhance embedding performance. Subsequently, the merging of general models and fine-tuned models was demonstrated to achieve superior zero-shot performance [256]. LM-Cocktail [257] pioneered testing on PLM-based embedders, weighing the parameters of the general embedder against those of the fine-tuned models on target tasks. The embedder with the weighted parameter can simultaneously preserve the target task’s performance while maintaining the general capabilities across other tasks. With the introduction of methods such as task arithmetic [258] and DARE [259], this technique was applied to model checkpoints fine-tuned for different tasks on the same starting point (identical pre-trained models). In the field of text embedding, different tasks do not necessarily reinforce each other; in fact, they can even conflict. For instance, Cpt [194] observes that retrieval/classification performance improved while STS performance declined as the number of training steps increased; InstructOR [67] observes that training symmetric and asymmetric tasks together without adding different instructions leads to a decline in performance. These issues were initially overlooked, then partially addressed by the methods based on instruction tuning. However, the data size across different tasks can vary dramatically. In traditional multi-task learning, this issue can only be addressed through technically challenging methods, such as task gradient weighting [260] or resampling [261]. As a post-processing technique, model merging is significantly less difficult to implement than multi-task learning, and its performance has progressively reached and even surpassed that of the latter during its development [262]. Therefore, some empirical studies have applied task-based arithmetic and its variants to general embedding [263] and dense retrieval [264, 265]. In the implementation of the advanced embedder, Lychee [205] divides all training data into four categories: (1) basic relevance retrieval, (2) code retrieval, (3) tool retrieval, and (4) complex instruction-based retrieval, fine-tuning one embedder for each category dataset from the same initiation point. Then, the merging method called spherical linear interpolation (SLERP) [^11] https://github.com/Digitous/LLM-SLERP-Merge is introduced to merge four embedders with an efficient hyper-parameter search. Note that Qwen3 Embedding [219] also utilizes SLERP, but the method for acquiring the merged checkpoints remains unclear. Recently, CoDiEmb [209] A first employed the model soups approach to fuse separate models for the IR and STS tasks. Subsequently, it applied finer-grained layer-wise weights to adjust the soup models of IR and STS, ultimately yielding the final checkpoint.
@@ 4.4 Commercial Service
Many commercial companies provide generic text embedding services to support traditional NLP tasks and RAG. In English text embedding service, it mainly includes OpenAI  [^12] https://openai.com/index/introducing-text-and-code-embeddings, Google  [^13] https://cloud.google.com/vertex-ai/generative-ai/docs/embeddings/get-text-embeddings, Amazon  [^14] https://docs.aws.amazon.com/bedrock/latest/userguide/titan-embedding-models.html, Alibaba Cloud [^15] https://www.alibabacloud.com/help/en/model-studio/developer-reference/general-text-embedding, ByteDance [^16] https://seed1-5-embedding.github.io/, Voyage  [^17] https://docs.voyageai.com/docs/embeddings, Cohere  [^18] https://cohere.com/blog/introducing-embed-v3 and Jina  [^19] https://jina.ai/embeddings. Due to the trade secrets involved, only part of companies have revealed some of the core technology behind their services. For example, OpenAI simultaneously provides text embeddings with different dimensional settings for different levels of time-storage sensitivity in practice, where matryoshka representation learning (MRL) [266] is considered to be the key to obtaining the embedding with flexible dimensions using only one encoder. Google demonstrates that Gecko [156], which achieves superior performance on 1B scale and 786 dimensions, relies heavily on a two-stage synthetic data generation process with LLMs: the first step generates diverse tasks and queries based on the instruction and the second step generates positive and negative samples based on the obtained tasks and queries. As shown in Table 4, many commercial companies publish their own technical reports. However, with the exception of a few that are fully open-source [211], most do not release their source code or data. Furthermore, they do not disclose all details regarding their data processing or training methods. Some companies (such as Bytedance and Voyage) have published results on public benchmarks, but only provided API services without further details. In addition, there have been the attempt [267] that attempt to distill knowledge through embeddings obtained from commercial APIs, successfully obtaining a local embedder with similar performance.